OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @ganeshie8

OpenStudy (anonymous):

\[Given~\log _{3}2=0.631~and~\log _{3}5=1.465,find~the~values~of:\]\[\log _{3}30\]

ganeshie8 (ganeshie8):

write \(30\) as \(3\times 2\times 5\) and use below log property \[\large \log_b (xyz) = \log_b( x) + \log_b( y) + \log_b (z)\]

OpenStudy (anonymous):

\[\log _{3}2+\log _{3}3+\log _{3}5\]

ganeshie8 (ganeshie8):

+1 for writing the enitre question in neat latex :)

OpenStudy (anonymous):

okay i get it. Answer is 3.096

ganeshie8 (ganeshie8):

\[\large \color{Red}{\checkmark}\]

OpenStudy (anonymous):

Thnx @ganeshie8 :)

ganeshie8 (ganeshie8):

yw:)